# Machine Learning for Detecting Steering in Qutrit-Pair States

### v2

@@ Abstract
keywords:
quantum steering, machine learning, qutrit-qutrit systems, quantum steering ellipsoid, quantum steerability bounds
@@
03.67.-a, 03.65.Ud, 03.65.-w
1 Introduction
[…]
a given quantum state is steering or not through numerical simulations [21].
Recently, machine learning has garnered significant interest for its integration with quantum information, particularly following its successful applications in entanglement discrimination [22, 23], non-locality detection [24] and quantum steering detection in two-qubit systems. In 2019, Ren et al. [25] first applied the support vector machine (SVM) algorithm for quantum steering of arbitrary two-qubit states. In 2020, Zhang et al. [26] used a back-propagation neural network to train quantum steering classifiers, and Zhang et al. (2021) [27] applied semi-supervised learning methods to this task. More recently, in 2024, Zhang et al. [28] explored steerability detection in qubit-qutrit systems using various machine learning techniques. However, machine learning for high-dimensional quantum steering detection remains limited. High-dimensional quantum steering [29] has gained attention for its enhanced noise resilience, but there are few related works in both theoretically predicted and experimentally confirmed [30].
In this paper, we focus on quantum steering detection for the qutrit-qutrit system by different machine learning methods. Notable contributions made in this paper are as follows.
  • We build a dataset for steerability detection utilizing SDP. For the full-information feature $F_{1}$, we demonstrate that artificial neural networks (ANNs) achieve high classification accuracy and generalization, outperforming support vector machines (SVMs).
  • We propose steering ellipsoid-like feature $F_{2}$, which significantly enhances the performance of each of our models.
  • We construct the first rigorously constructed, accurately labeled dataset based on theoretical foundations. This dataset enables machine learning models to exhibit outstanding accuracy and generalization capabilities, independent of the.choice of features.
  • As applications, we utilize these models to investigate the steerability boundaries of isotropic states and partially entangled states, and find new steerable states.
The rest of the paper is organized as follows. In the next section, we introduce basic concepts and SDP detecting method of quantum steering needed in later sections. In Sec.3, we pioneer the construction of SDP-labeled and accurately labeled datasets for qutrit-qutrit systems. Compared with the full information feature $F_{1}$, our steering ellipsoid-like feature $F_{2}$ can train SVM models with higher classification accuracy and more robust generalisation. In the same sense, accurately labeled dataset performs better than SDP-labeled dataset. In Sec.4 and Sec.5, we train ANN and ensemble learning classifiers. Though analysis, we find that ANN models achieve better generalization performance than SVM models for the feature $F_{1}$, and ensemble learning models perform best for the feature $F_{2}$. Sec.6 is devoted to the steerability boundaries of isotropic states and partially entangled states, and find new steerable states by using above trained model. Finally, we summarize our results in Sec.7.
2 PRELIMINARIES
@@ 3 Detecting the steerability by SVM
3.1 Data production with feature $F_{1}$
3.1.1 Data production
@@
    $S_{x}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}0&1&0\\ 1&0&1\\ 0&1&0\\ \end{array}\right),S_{y}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}0&-i&0\\ i&0&-i\\ 0&i&0\\ \end{array}\right),S_{z}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}1&0&0\\ 0&0&0\\ 0&0&-1\\ \end{array}\right)$ (14)
@@ 3.1.2 Training and testing
The fig. 1 shows that the classification accuracy of SVM models on random states can be kept over $85\%$ (See appendix table 4 for detailed data) except $m=4$. In the case $m=4$, the classification accuracy can be improved to 89.4% on train set and the test accuracy improved to 87.5% by regenerating 20000 random states with 10000 positive and 10000 negative samples. Consistent with our intuition, increasing the amount of data can improve the classification accuracy of the model.
@@
We uniformly select the parameter $\eta$ within intervals $[0,\frac{H_{3}-1}{2}]$ and $[\frac{H_{3}-1}{2},1]$ to generate 2000 examples of unsteering isotropic states and 2000 examples of steering isotropic states, respectively. Then we apply the above pre-trained SVM models to classify the dataset of isotropic states. However, the classification accuracy remains at $50\%$, indicating a lack of robust generalization performance. Compared with the standard deviation heatmaps of the training set data, characterized by $F_{1}$ for the random state $(m=3)$, the corresponding heatmaps for isotropic states has more locations containing zeros. See Figure. 2. The reason for the lack of generalisation ability of our SVM models may be that the test set generated by isotropic states is in the margin of random states set.
@@
3.2 Data with feature $F_{2}$ production
@@ 3.2.1 Data production
3.2.2 Training and testing
Similar to the research in Sec.3.1.2, we can train SVM models for feature $F_{2}$, and obtain the accuracy of these well-trained SVM models in the test set. It is encouraging that they not only slightly improve the classification accuracy on random states, but also have robust generalization performance, i.e., they can classify the dataset of isotropic states with the accuracy over $89.7\%$ (The details of the data are in table 5).
@@ 3.2.2 Training and testing
3.3 Data production with accurate labels
3.3.1 Data production
@@ 3.3.2 Training and testing
Based on the results in table 1, it is evident that the well-trained SVM models with accurately labeled datasets, in contrast to the original datasets, yields significantly better performance. Specifically, the accurately labeled datasets outperform the originals in terms of cross-validation accuracy, testing accuracy, and generalization capability. These findings strongly indicate that the use of accurate labels contributes to reduced quantum steering classification errors and enhances the overall effectiveness of SVM model training.
@@
      partially-entangle       94.90%       91.90%
4 Detecting the steerability by the artificial neural network
@@ 4.1 Training and testing
Exchange the feature $F_{1}$ for the feature $F_{2}$, the classification accuracy of ANN models is improved on random states. The fig. 6 shows that the accuracy of ANN models prediction with feature $F_{2}$ is improved to around 87% on random states (The details of the data are in table 7).
@@
Whereas the excellent performance of the SVM models on accurately labeled datasets, we use the same high-quality datasets to train and test ANN models. As shown in table 2, the ANN models perform very well on the accurately labeled datasets, exhibiting not only high accuracy but also robust generalization capabilities. Notably, when comparing the datasets with the accurate labels for feature $F^{\prime}_{2}$, the ANN models achieve higher cross-validation accuracy, testing accuracy, and overall generalization ability on the datasets with feature $F^{\prime}_{1}$.
@@
5 Detecting the steerability by ensemble learning
@@
5.1 Training and testing
When we use the datasets with feature $F_{1}$, no decision tree can classify successfully. Exchange the feature $F_{1}$ for the feature $F_{2}$, the classification accuracy of ensemble learning models is the highest compared with SVM and ANN models. The fig. 8 shows that ensemble learning models achieve an accuracy of around 87.5% for random states steering classification and over 91% for isotropic states classification (The details of the data are in table 8).
@@ 5.1 Training and testing
6 Detection of quantum steering bounds
@@ 7 Conclusion
[…]
This research was partially supported by the National Natural Science Foundations of China (Grant No. 11901317), the China Postdoctoral Science Foundation (Grant No. 2020M680480), the Fundamental Research Funds for the Central Universities (Grant No. 2023MS078), and the Beijing Natural Science Foundation (Grant No. 1232021).
P.W. and Z.Y.L. contributed equally to this work.
[…]
The authors declare that they have no conflict of interest.
Appendix Appendix

### v3

@@ Abstract
preprint: APS/123-QED
I Introduction
@@
Recently, machine learning has garnered significant interest for its integration with quantum information, particularly following its successful applications in entanglement discrimination [22, 23], non-locality detection [24] and quantum steering detection in two-qubit systems. In 2019, Ren et al. [25] first applied the support vector machine (SVM) algorithm for quantum steering of arbitrary two-qubit states. In 2020, Zhang et al. [26] used a back-propagation neural network to train quantum steering classifiers, and Zhang et al. (2021) [27] applied semi-supervised learning methods to this task. More recently, in 2024, Zhang et al. [28] explored steerability detection in qubit-qutrit systems using various machine learning techniques. However, machine learning for high-dimensional quantum steering detection remains limited. High-dimensional quantum steering [29] has gained attention for its enhanced noise resilience, but there are few related works in both theoretical predictions and experimental confirmations [30]. In this paper, we extend quantum steering detection to the qutrit-qutrit system, applying machine learning methods to address this challenge.
In this paper, we will detect quantum steering of qutrit-qutrit states via different machine learning techniques. In Sec.II, we introduce basic concepts and SDP detecting method of quantum steering. In Sec.III, we pioneer the construction of SDP-labeled and accurately labeled datasets for qutrit-qutrit systems. Compared with the full information feature $F_{1}$, our steering ellipsoid-like feature $F_{2}$ can train SVM models with higher classification accuracy and more robust generalisation. In the same sense, accurately labeled dataset performs better than SDP-labeled dataset. In Sec.IV and Sec.V, we train ANN and ensemble learning classifiers. Though analysis, we find that ANN models achieve better generalization performance than SVM models for the feature $F_{1}$, and ensemble learning models perform best for the feature $F_{2}$. In Sec.VI, we explore the above trained model to investigate the steerability boundaries of isotropic states and partially entangled states, and find new steerable states. Finally, we summarize our results in Sec.VII.
II PRELIMINARIES
@@ III Detecting the steerability by SVM
III.1 Data production with feature $F_{1}$
III.1.1 Data production
@@
    $S_{x}$ $=$ $\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}0&1&0\\ 1&0&1\\ 0&1&0\\ \end{array}\right),S_{y}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}0&-i&0\\ i&0&-i\\ 0&i&0\\ \end{array}\right),$ (11)
    $S_{z}$ $=$ $\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}1&0&0\\ 0&0&0\\ 0&0&-1\\ \end{array}\right)$ (15)
@@ III.1.2 Training and testing
The Fig. 1 shows that the classification accuracy of SVM models on random states can be kept over $85\%$ (See appendix 4 for detailed data) except $m=4$. In the case $m=4$, the classification accuracy can be improved to 89.4% on train set and the test accuracy improved to 87.5% by regenerating 20000 random states with 10000 positive and 10000 negative samples. Consistent with our intuition, increasing the amount of data can improve the classification accuracy of the model.
@@
We uniformly select the parameter $\eta$ within intervals $[0,\frac{H_{3}-1}{2}]$ and $[\frac{H_{3}-1}{2},1]$ to generate 2000 examples of unsteering isotropic states and 2000 examples of steering isotropic states, respectively. Then we apply the above pre-trained SVM models to classify the dataset of isotropic states. However, the classification accuracy remains at $50\%$, indicating a lack of robust generalization performance. Compared with the standard deviation heatmaps of the training set data, characterized by $F_{1}$ for the random state $(m=3)$, the corresponding heatmaps for isotropic states has more locations containing zeros. See Fig. 2. The reason for the lack of generalisation ability of our SVM models may be that the test set generated by isotropic states is in the margin of random states set.
@@
III.2 Data with feature $F_{2}$ production
@@ III.2.1 Data production
III.2.2 Training and testing
Similar to the research in Sec.3.1.2, we can train SVM models for feature $F_{2}$, and obtain the accuracy of these well-trained SVM models in the test set. It is encouraging that they not only slightly improve the classification accuracy on random states, but also have robust generalization performance, i.e., they can classify the dataset of isotropic states with the accuracy over $89.7\%$ (The details of the data are in Tab. 5).
@@ III.2.2 Training and testing
III.3 Data production with accurate labels
III.3.1 Data production
@@ III.3.2 Training and testing
Based on the results in Tab. 1, it is evident that the well-trained SVM models with accurately labeled datasets, in contrast to the original datasets, yields significantly better performance. Specifically, the accurately labeled datasets outperform the originals in terms of cross-validation accuracy, testing accuracy, and generalization capability. These findings strongly indicate that the use of accurate labels contributes to reduced quantum steering classification errors and enhances the overall effectiveness of SVM model training.
@@
IV Detecting the steerability by the artificial neural network
@@ IV.1 Training and testing
Exchange the feature $F_{1}$ for the feature $F_{2}$, the classification accuracy of ANN models is improved on random states. The Fig. 6 shows that the accuracy of ANN models prediction with feature $F_{2}$ is improved to around 87% on random states (The details of the data are in Tab. 7).
@@
Whereas the excellent performance of the SVM models on accurately labeled datasets, we use the same high-quality datasets to train and test ANN models. As shown in Tab. 2, the ANN models perform very well on the accurately labeled datasets, exhibiting not only high accuracy but also robust generalization capabilities. Notably, when comparing the datasets with the accurate labels for feature $F^{\prime}_{2}$, the ANN models achieve higher cross-validation accuracy, testing accuracy, and overall generalization ability on the datasets with feature $F^{\prime}_{1}$.
@@
V Detecting the steerability by ensemble learning
@@
V.1 Training and testing
When we use the datasets with feature $F_{1}$, no decision tree can classify successfully. Exchange the feature $F_{1}$ for the feature $F_{2}$, the classification accuracy of ensemble learning models is the highest compared with SVM and ANN models. The Fig. 8 shows that ensemble learning models achieve an accuracy of around 87.5% for random states steering classification and over 91% for isotropic states classification (The details of the data are in Tab. 8).
@@ V.1 Training and testing
VI Detection of quantum steering bounds
@@ VII Conclusion
Acknowledgements.
This research was partially supported by the National Natural Science Foundations of China (Grant No. 11901317), the China Postdoctoral Science Foundation (Grant No. 2020M680480), the Fundamental Research Funds for the Central Universities (Grant No. 2023MS078), and the Beijing Natural Science Foundation (Grant No. 1232021). P.W. and Z.Y.L. contributed equally to this work.
Appendix A Appendixes